Do the following problems using your financial calculator. Show all your work and complete the register.

  1. In 2013, Barrick Gold, a Canadian mining company, adopted a plan to accumulate a remediation fund beginning Aug. 2, 2018 at an estimated cost of $32 million. Barrick plans to make equal semiannual payments into a fund earning 5.2% interested compounded semiannually. The first deposit is scheduled for Aug. 1, 2013. Determine the required semiannual deposit. First determine how many payments Barrick will be making. Assume the last payment is made on Feb. 1, 2018 but the funds are not needed until Aug. 1, 2018. (5 points)

(((((((((((((((((((((    I want to make the first question clear.

The plan was adopted in 2013 (not 2015, typo) so that there would be $32 million in a fund by Aug. 1, 2018.The last payment is made on Feb. 1, 2018 but the funds are not withdrawn until Aug. 1, 2018. Count the number of payments that will be made (make sure you understand the total payments will sit for six months in the bank before being withdrawn).

So to be clear. The first payment is made on Feb. 1, 2013. A new payment is made every 6 months until Feb. 1, 2018. There needs to be 32 million in the account on Aug. 1, 2018. It is a sinking fund))))))))))))))))))))))))))

N =            I/Y =               PV =                PMT =             FV =

                             

  1. Deborah Gianoulis won $6.5 million in the Georgia Lottery. She must choose how she wants her prize to be paid. First, Deborah can elect to receive 26 annual payments, with the first payment due immediately. Second, she can elect to receive a single payment for the entire amount. However, if the single payment option, the winning prize is reduced to 60% of the winning. What rate would make Deborah indifferent about these two options? (10 points)

26 Annual Payments

N =            I/Y =               PV =                PMT =             FV =

1 Payment

Amount:
Answer #1

1.

For Periodic Deposit
p = (A x i)/ ( 1+1)^n-1
n = 11
A = 32 Million
I = 5.2/2 = 2.6%
=32000000*.026/((1.026)^11-1)
2550297.283
N 11
i/y ( semi annually) 2.60%
PMT (in $)      25,50,297.28
FV (in $) 3,20,00,000.00
PV = (PMT X (1-(1+i)^-n)) / i
PV = (2550297.28 X (1-(1+.026)^-11)) /.026 2,41,28,426.97
PV (in $) 2,41,28,426.97

2.

Prize Amount (in$)        65,00,000
Lumpsum payout (in $)        39,00,000
PV = FV/(1+i)^n
n = 26
PV = 3900000
FV = 6500000
I = (FV/PV)^1/t -1
=(6500000/3900000)^1/26 - 1
i = 1.98%